OpenStudy (anonymous):

solve for x: 5 x ---- = ---- 3x+4 x+2

OpenStudy (compassionate):

First, in order to solve this, you want to see if you can get any factors.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x(3x + 4) = 5(x + 2) 3x^2 + 4x = 5x + 10 3x^2 - x - 10 (3x + 5)(x + 6)

OpenStudy (anonymous):

x = -5/3 and -6